I am doing permutation problems using the formula nPr = n!/(n-r)!. The problem I am having difficulty with is 6Pn = 30. I can get it this far:
6! / (6-X)! = 30
6! / 30 = (6-X)!
24 = (6-X)! Then I can't figure out how to get X. Please help. Thanks.

Answer by stanbon(29602)

6! / (6-X)! = 30
6! / 30 = (6-X)!
-------------------
720/30 = (6-x)!
24 = (6-x)!
4! = (6-x)!
6-x = 4
x = 2
==========
Cheers,
Stan H.